# Neuro Symbolic Knowledge Reasoning for Procedural Video Question Answering with Knowledge Module Learning

### v3

@@ Abstract
We introduce PKR-QA (Procedural Knowledge Reasoning Question Answering), a new benchmark for question answering over procedural tasks that require structured reasoning. PKR-QA is constructed semi-automatically using a procedural knowledge graph (PKG), which encodes task-specific knowledge across diverse domains. The PKG is built by curating and linking information from the COIN instructional video dataset and the ontology, enriched with commonsense knowledge from ConceptNet and structured outputs from Large Language Models (LLMs), followed by manual verification. To generate question-answer pairs, we design graph traversal templates where each template is applied systematically over PKG. To enable interpretable reasoning, we propose a neurosymbolic approach called Knowledge Module Learning (KML), which learns procedural relations via neural modules and composes them for structured reasoning with LLMs. Experiments demonstrate that this paradigm improves reasoning performance on PKR-QA and enables step-by-step reasoning traces that facilitate interpretability. Code and dataset will be released soon https://github.com/LUNAProject22/KML.
1 Introduction
Interest in understanding procedural tasks is growing, driven by applications in cooking, machinery repair, medical procedures, and daily activities [4, 66, 35], as reflected by platforms like WikiHow. Such tasks consist of sequences of steps, where each step serves a specific purpose and tools are used accordingly. Human understanding of these tasks arises from lifelong learning [49], enriched by commonsense knowledge of objects and their affordances, as well as the ability to reason about temporal and causal dependencies within multi-step processes. This enables individuals to infer not just what to do, but why and how each step contributes to the overarching goal. Emulating this procedural knowledge reasoning—a key aspect of human cognition—is essential for machines to assist in complex real-world tasks.
Despite the rise of visual question answering and general knowledge-based benchmarks, there remains a gap in evaluating a model’s ability to perform procedural knowledge reasoning. To address this, we introduce PKR-QA (Procedural Knowledge Reasoning Question Answering), a new benchmark designed to assess Vision Language Models (VLMs) and Neurosymbolic (NS) methods on diverse procedural reasoning capabilities, including multi-hop, deductive, probabilistic, contextual, causal, and counterfactual reasoning. Many procedural tasks require domain-specific knowledge and an understanding of task structures, which can be effectively captured using knowledge graphs. To construct the PKR-QA questions and answers, we build a procedural knowledge graph (PKG) that encodes temporal relationships and causal links among key concepts. Our dataset challenges models to answer procedural questions by combining information from the video with external knowledge about the task. Unlike previous video QA benchmarks that primarily test visual comprehension, i.e., reasoning over what is in the video, our dataset emphasizes task-centric reasoning and knowledge grounded in the procedure. This introduces new challenges that not only demands visual understanding and reasoning, but also to reason about the procedural knowledge of the task [54, 62, 25, 56, 50].
While VLMs demonstrate impressive reasoning and knowledge capabilities, their internal reasoning mechanism is not that transparent. Techniques like chain-of-thought, graph-of-thought, and tree-of-thought aim to enhance VLM reasoning [53, 6, 58], yet the reasoning remains implicit due to the lack of constraints on intermediate variables. Moreover, reasoning and execution are often entangled within the VLM’s internal mechanism. To address this, methods like ViperGPT [47] decouple reasoning from execution by generating executable programs. Inspired by such approaches and neurosymbolic (NS) models [22, 32, 3, 38, 18, 19, 10, 12], we propose Knowledge Module Learning (KML), which models knowledge-based relations as parameterized, trainable modules leveraging on the modern embeddings. KML generate the program of knowledge modules to answer knowledge-based reasoning questions as compositions of knowledge modules using the PKG schema and LLMs. KML explicitly answer knowledge-based reasoning questions leveraging the powerful reasoning mechanism of moderns LLMs allowing it to execute reasoning steps in a more transparent manner to derive the answer. KML supports explicit knowledge-based reasoning, constrained to a well-defined set of knowledge operations, plays a crucial role beyond procedural tasks in domains that require structured decision-making, reliability, and interpretability. KML provides explicit reasoning steps that can be systematically verified, debugged, and refined. Unlike end-to-end black-box models, which often lack transparency in their decision-making, structured reasoning through predefined operations enables more interpretable, trustworthy, and controllable AI systems. Our dataset and the approach promote these aspects.
Our contributions are three-fold. First, we introduce PKR-QA, a new benchmark to test procedural knowledge reasoning by question answering. Second, we propose KML, a neuro-symbolic method that executes LLM-generated programs over knowledge graphs using relation-specific neural modules. Third, we benchmark PKR-QA with state-of-the-art VLMs and NS methods. KML variants outperform all baselines, highlighting the benefit of structured knowledge for interpretable and controllable procedural reasoning. We believe PKR-QA will serve as a valuable resource to advance trustworthy AI in this emerging domain.
2 Related Work
Understanding procedural and instructional videos remains a significant challenge in computer vision, as explored in prior research [33, 4, 66, 64, 35, 8, 45]. The authors in [4] introduced the key step recognition and procedure-aware video representation learning approach using step transitions is presented in [66]. We leverage on these works.
@@ 2 Related Work
Knowledge-based Visual Question Answering has become popular in recent years [31, 9, 36, 42]. The majority of the past benchmarks on knowledge-based visual question answering are for images while more recently authors in  [50] presented a benchmark to evaluate situated and open-world common-sense reasoning in videos with compositionality, temporality, and causality. While the motivation of our work shares similarities with [50], there are notable distinctions. Their focus lies in situated commonsense reasoning grounded in the specific contexts depicted in videos, whereas we concentrate on testing the procedural understanding of models given some partial information such as a step of a task.
Integrating External Knowledge with Video Understanding has been explored through knowledge graphs (KGs) for tasks like activity recognition [29, 14] and visual commonsense reasoning [26]. VidSitu [41] links events to semantic roles, while TVQA+[23] uses script knowledge for story-based QA. However, these works focus on descriptive reasoning rather than procedural reasoning.
@@
Large VLMs (e.g., Flamingo [2], Video-ChatGPT [30]) and neurosymbolic frameworks like ViperGPT [47], MoreVqa [34] and [11] decouple reasoning (e.g., program generation) from execution (e.g., API calls). While ViperGPT uses predefined functions for visual queries, it lacks mechanisms to learn domain-specific knowledge modules or constrain reasoning to procedural logic. Similarly, chain-of-thought prompting [53] improves transparency in LLMs/VLMs but remains dependent on the model’s internal knowledge, which may be unreliable for specialized domains. Our framework advances this by (1) training lightweight, interpretable Knowledge Modules directly from a domain specific KG and (2) constraining reasoning to predefined operations ensuring alignment with domain knowledge.
@@
Our work is also related to NLP research such as [1, 57, 65, 43], which employ logical queries to extract information from a KG. This emphasis on curated, domain-specific knowledge is especially valuable for AI assistants operating in contexts where domain expertise, rather than general commonsense reasoning, is crucial for success. Prior knowledge completion works that learns embeddings of entities and relations in KGs are also related to us [7, 52]. In contrast, KML leverage on modern embeddings to learn relational mappings as neural functions, and therefore fundamentally different.
3 PKG Construction and Dataset Creation
@@
Motivated by recent advances in the semi-automated construction of knowledge-based question answering datasets [16], we introduce the Procedural Knowledge Reasoning Question Answering (PKR-QA) dataset. It is built on a Procedural Knowledge Graph (PKG) that integrates information from the COIN training set, the COIN ontology, GPT-4o-generated annotations, and external commonsense knowledge from ConceptNet, followed by human verification. To enable systematic QA generation, we define a set of question templates and associated Cypher queries [13], which are executed over the knowledge graph to retrieve correct answers. In the following sections, we detail the construction of both the knowledge graph and the QA dataset.
3.0.1 Defining PKG’s Schema (PKGS).
@@
A knowledge graph schema provides a high-level abstraction of the graph, specifying the types of entities that exist and the valid relationships that can occur between them. It serves as a blueprint for structuring the data and interpreting the semantics of the graph. In our case, PKGS contains nodes where each node correspond to a entity type ($\mathcal{E}$) and edges represent relation types ($\mathcal{R}$) (Figure 1). The core entity types in PKGS include Domain, Task, Step, Action, Object, Tool, and Purpose. Relation types capture meaningful procedural links, such as task-step associations or tool usage (e.g. HAS_TOOL), and may carry attributes like id, type, or additional semantic metadata.
3.0.2 Populating PKG.
@@
Based on this schema, we instantiate PKG by populating it with specific entity and relation instances. Each entity $e\in\mathcal{E}$ has three main attributes: type, name, and id. Relations $r\in\mathcal{R}$ represent specific connections between entity instances. To construct this graph, we integrate annotations from the COIN dataset [48], fine-grained information extracted using GPT-4o [20], and commonsense knowledge from ConceptNet [44].
@@
COIN-based Data (Domain, Task, Step): We extract the procedural structure from the training split of the COIN dataset, which provides annotations for domains, tasks, and steps. For each unique instance of these entities, we create corresponding nodes in the graph. We define HAS_TASK edges to link each domain to its associated tasks, and HAS_STEP edges to connect tasks to their constituent steps. To capture procedural flow, we analyze step sequences from all training videos and construct HAS_NEXT_STEP relations that represent the temporal ordering of steps. Additionally, we introduce special START and END nodes to explicitly mark task boundaries. Each HAS_NEXT_STEP edge is annotated with its observed frequency to model empirical transition likelihoods between steps.
@@
LLM-Augmented Data (Action, Object, Tool): To enhance procedural specificity, we use GPT-4o to extract action-object pairs from step descriptions. For instance, the phrase “remove the tire” yields the action “remove” and the object “tire”. Because COIN lacks explicit tool annotations, we further prompt GPT-4o to infer potential tools for each step. The tool lists are manually verified. We standardize terms (e.g., “scissor” $\rightarrow$ “scissors”), consolidate duplicates (e.g., “marker”, “marker pen”), and unify task-specific variants (e.g., “toilet detergent” $\rightarrow$ “detergent”). Some task-sensitive distinctions are preserved (e.g., “filter” $\rightarrow$ “coffee filter”) when relevant to the procedural context.
@@
ConceptNet-based Data (Purpose): In procedural tasks, actions are typically performed—and tools or objects used—with specific purposes. To represent such intent in our graph, we incorporate commonsense knowledge from ConceptNet, a large-scale semantic network that connects words and phrases through meaningful, human-readable relations such as UsedFor, CapableOf, and IsA. Specifically, we extract potential purposes of actions, tools, and objects using the UsedFor and CapableOf relations. To contextualize these purposes within specific procedural steps and tasks, we use GPT-4o to infer task- and step-specific interpretations for each entity’s purpose (see the prompt template in Figure 7, Supplementary). This step ensures that the resulting knowledge is not generic but grounded in the procedural context in which the entity is used.
@@
To reduce redundancy and merge semantically similar purposes, we compute pairwise cosine similarities of Sentence-BERT embeddings [40] and apply a similarity threshold of 0.8, validated through manual inspection. Pairs of highly similar purposes are linked using a HAS_SIMILAR_PURPOSE edge. The resulting purpose-augmented graph is stored in Neo4j [^1] https://neo4j.com/, supporting efficient querying and reasoning via Cypher. All contents are manually verified to ensure the reliability of the KG. The constructed KG contains 2,954 unique entities and 12,484 relations. Detailed distributions of entities and relations are shown in Figure 5 (Supplementary).

### v5

@@ Abstract
In this work we present Knowledge Module Learning (KML) to understand and reason over procedural tasks that requires models to learn structured and compositional procedural knowledge. KML is a neurosymbolic framework that learns relation categories within a knowledge graph as neural knowledge modules and composes them into executable reasoning programs generated by large language models (LLMs). Each module encodes a specific procedural relation capturing how each entity type such as tools are related to steps, purpose of each tool, and steps of each task. Given a question conditioned on a task shown in a video, then KML performs multistep reasoning with transparent, traceable intermediate states. Our theoretical analysis demonstrated two desired properties of KML. KML satisfy strong optimal conditions for modelling KG relations as neural mappings, providing strong foundations for generalizable procedural reasoning. It also shows a bound on the expected error when it performs multistep reasoning. To evaluate this model, we construct a large procedural knowledge graph (PKG) consisting of diverse instructional domains by integrating the COIN instructional video dataset, and COIN ontology, commonsense relations from ConceptNet, and structured extractions from LLMs, followed by expert verification. We then generate question and answer pairs by applying graph-traversal templates over the PKG, constructing the PKR-QA benchmark for procedural knowledge reasoning. Experiments show that KML improves structured reasoning performance while providing interpretable step-by-step traces, outperforming LLM-only and black-box neural baselines. Code is publicly available at https://github.com/LUNAProject22/KML.
I Introduction
The procedural knowledge is the ability to understand and reason through how, why, and what sequence of steps are executed to accomplish a task. Procedural knowledge is fundamentally structured, sequential, and goal-oriented. It demands the handling of temporal dependencies and causal relationships in each step, while accounting for tool affordances and the functional intent behind every step. In humans, this knowledge is developed via lifelong experience and hierarchical abstraction. This process facilitates high-level generalization, allowing individuals to adapt to environmental constraints (such as missing instrumentation), predict future states (such as mental time travel [suddendorf1997mental]), and provide explainable rationales for their actions [thoe2022developing]. Replicating these sophisticated reasoning capabilities is critical for the advancement of autonomous systems. For embodied AI to achieve reliability in highly complex environments including household robotics, industrial maintenance, and clinical healthcare, it is well understood that they must transit from pattern recognition to a deep understanding of the procedural logic that can supports complex human tasks [ashutosh2024video, zhou2023procedure, nagasinghe2024not].
Recent advances in Vision–Language Models (VLMs) have demonstrated impressive performance on a wide range of multimodal reasoning tasks, including video understanding [li2024mvbench] and question answering [lei2018tvqa]. However, despite their strong pattern recognition and linguistic reasoning abilities, current VLMs largely rely on implicit, unstructured reasoning encoded within large neural networks. As a result, their reasoning processes are difficult to interpret, hard to constrain to domain-specific logic, and prone to hallucination when faced with incomplete or out-of-distribution procedural knowledge. This limitation becomes particularly critical in procedural domains, where correctness, reliability, and traceability of reasoning are as important as final accuracy.
Several recent efforts attempt to improve the transparency of LLM/VLM reasoning through prompting strategies such as chain-of-thought, tree-of-thought, and graph-of-thought reasoning [wei2022chain, besta2024graph, yao2023tree]. While these approaches encourage models to externalize intermediate reasoning steps, the underlying reasoning variables remain unconstrained and are not grounded in explicit symbolic structures or explicit logical reasoning. Consequently, reasoning and execution are still tightly coupled within the model, limiting the ability to verify, debug, or systematically improve reasoning behaviour, especially when domain-specific procedural constraints must be respected. Consequently, hallucinations and under explained results are still potential to occur.
To address this, methods like ViperGPT [suris2023vipergpt] decouple reasoning from execution by generating executable programs. In parallel, neuro-symbolic (NS) approaches aim to combine the flexibility of neural representations with the structure and interpretability of symbolic reasoning. However, many existing NS systems either rely on manually defined symbolic operators, which do not scale well to large procedural domains, or on fixed knowledge graph embeddings that struggle with compositional multi-hop reasoning and uncertainty propagation. Moreover, most prior NS approaches have focused on static scenes or short-horizon reasoning, leaving procedural, multi-step, temporally grounded reasoning in videos under explored [johnson2017inferring, mascharka2018transparency, andreas2016learning, perez2018film, hudson2019learning, hudson2018compositional, chen2021meta, endo2023motion, li2025imore].
To address these challenges, we propose Knowledge Module Learning (KML), a neuro-symbolic framework that explicitly models procedural relations as learnable neural modules and composes them into executable reasoning programs generated by large language models. Unlike black-box VLM reasoning, KML separates what reasoning steps should be performed (program synthesis) from how each step is executed (relation-specific neural knowledge modules). Each module corresponds to a well-defined relation in a procedural knowledge graph (PKG) and is trained to map between entity types (e.g., steps to tools, tools to purposes) using contrastive learning.
This design yields several key advantages. First, reasoning is explicit and traceable, producing interpretable intermediate states that correspond to meaningful procedural entities. Second, reasoning is constrained by the schema of the PKG, ensuring alignment with domain knowledge while still allowing flexibility through neural generalization. Third, KML naturally supports uncertainty-aware multi-hop reasoning, enabling it to handle ambiguous grounding and incomplete observations—common challenges in real-world video understanding.
To rigorously evaluate procedural knowledge reasoning, we introduce PKR-QA, a large-scale benchmark constructed from a curated procedural knowledge graph integrating instructional video annotations, common-sense knowledge, and LLM-assisted extraction followed by human verification. Unlike existing video QA benchmarks that emphasize visual recognition or situational common-sense [wu2024star, yu2019activitynet, li2024mvbench, xu2017video, wang2024sok], PKR-QA focuses on task-centric procedural reasoning, requiring models to infer purposes, anticipate steps, reason over alternatives, and generalize across tasks.
Finally, beyond empirical performance, we provide a theoretical analysis of KML, establishing sufficient conditions under which neural knowledge modules learn relation mappings that separate valid from invalid entities and deriving bounds on reasoning error under multi-step composition. This theoretical grounding, combined with extensive ablations and qualitative analyses, positions KML as a principled and scalable approach to trustworthy procedural reasoning.
This is an extended version of our AAAI 2026 paper [nguyen2026pkr]. We have included additional theoretical analysis of KML which allows us to better understand the capabilities of KML. Further, we have included additional ablation studies showing the impact of each component in our VQA architecture. We also extended the KML to perform logic-based operators such as logical AND OR and NOT and evaluate the performance.
II Related Work
@@
II-A Procedural and Instructional Video Understanding
Understanding instructional and procedural videos has been an active area of research in computer vision, with early works focusing on representation learning from narrated videos and step recognition [miech2020end, ashutosh2024video, zhou2023procedure, zhong2023learning, nagasinghe2024not, chang2020procedure, sun2022plate]. More recently tasks such as goal-inference [ee2025deduce, wan2025infer, roy2024predicting] for procedural tasks in both robotics and vision are getting popular. Datasets such as COIN [tang2019coin] and ActivityNet-QA [yu2019activitynet] have enabled progress in recognizing steps, actions, and temporal structure in instructional content. Subsequent works introduced procedure-aware representations and planning-oriented models that explicitly model step transitions, task graphs, and long-horizon dependencies. While these approaches significantly advance procedural perception, they primarily focus on recognizing or predicting steps rather than reasoning over explicit procedural knowledge such as tool affordances, purposes, or cross-task generalization. In contrast, our work targets procedural knowledge reasoning, where models must integrate observed video evidence with external structured knowledge to answer complex, multi-hop questions.
II-B Knowledge-based Visual Question Answering
@@
Knowledge-based visual question answering has become popular in recent years [marino2019ok, chang2022webqa, park2020visualcomet, schwenk2022okvqa]. The majority of the past benchmarks on knowledge-based visual question answering are for images while more recently authors in  [wang2024sok] presented a benchmark to evaluate situated and open-world common-sense reasoning in videos with compositionality, temporality, and causality. While the motivation of our work shares similarities with [wang2024sok], there are notable distinctions. Their focus lies in situated commonsense reasoning grounded in the specific contexts depicted in videos, whereas we concentrate on testing the procedural understanding of models given some partial information such as a step of a task.
II-C Neuro-Symbolic and Modular Reasoning
Neuro-symbolic reasoning frameworks aim to bridge neural perception with symbolic structure [de2020statistical, abbe2022learning]. Modular networks, neural state machines [hudson2019learning], and program-based reasoning models enable compositional reasoning by assembling neural components according to predicted programs [johnson2017inferring, mascharka2018transparency, andreas2016learning, perez2018film, hudson2019learning, hudson2018compositional, chen2021meta, endo2023motion, li2025imore]. More recent systems, such as ViperGPT [suris2023vipergpt] and related approaches, leverage LLMs to generate executable programs that invoke visual or symbolic functions. Despite their promise, existing modular systems typically rely on predefined, hand-engineered functions or operate over limited sets of symbolic operators. They also lack mechanisms to learn domain-specific relational knowledge directly from data. In contrast, KML learns relation-specific neural modules directly from a procedural knowledge graph, allowing it to scale to rich procedural domains while maintaining interpretability.
II-D Knowledge Graph Embeddings and Reasoning
Knowledge graph embedding methods such as TransE [bordes2013translating], TransH [wang2014knowledge], and RotatE [sun2019rotate] represent entities and relations in continuous vector spaces and support link prediction and multi-hop reasoning through embedding composition. Mostly, these methods only learn the embeddings using some geometric properties and assumptions of the vector spaces that they operate. In contrast KML learns neural networks to represent relations as mapping functions, and entities as embeddings as that operates through relations as mapping functions. The closest idea to our work is the pioneering work of NTN [socher2013reasoning]. However, NTN is only able to predict a confidence score for a given triplet and multi-hop reasoning becomes challenging with it’s use of complex bilinear neural network models to encode relations. Furthermore, these early methods are effective for knowledge completion, yet they are not designed to integrate perceptual grounding or to provide interpretable intermediate reasoning steps.
KML differs fundamentally from KG embedding approaches by treating relations as neural mapping functions rather than algebraic operators, enabling uncertainty-aware reasoning and seamless integration with multimodal grounding. Furthermore, KML’s execution traces align with symbolic reasoning paths, offering interpretability that embedding-based approaches lack.
II-E LLM-Based Reasoning and Program Synthesis
LLMs have demonstrated strong reasoning capabilities through prompting strategies such as chain-of-thought [wei2022chain], tree-of-thought [yao2023tree], and graph-of-thought [besta2024graph]. These methods improve reasoning transparency but do not enforce semantic constraints on intermediate steps, often leading to hallucinations or inconsistent reasoning in structured domains.
@@ II-E LLM-Based Reasoning and Program Synthesis
Large VLMs such as Flamingo [alayrac2022flamingo], Video-ChatGPT [maaz2023video]) and neurosymbolic frameworks like ViperGPT [suris2023vipergpt], MoreVqa [min2024morevqa] and [choudhury2024video] decouple reasoning (e.g., program generation) from execution (e.g., API calls). While ViperGPT uses predefined functions for visual queries, it lacks mechanisms to learn domain-specific knowledge modules or constrain reasoning to procedural logic.
KML leverages LLMs for program synthesis, but constrains execution to a predefined set of learned knowledge modules aligned with a procedural schema. This decoupling allows KML to benefit from the reasoning flexibility of LLMs while ensuring correctness, interpretability, and alignment with domain knowledge.
II-F Integrating External Knowledge with Video Understanding
External Knowledge has been explored through knowledge graphs (KGs) for tasks like activity recognition [ma2022visual, ghosh2020all] and visual commonsense reasoning [lin2019kagnet]. VidSitu [sadhu2021visual] links events to semantic roles, while TVQA+[lei2019tvqa+] uses script knowledge for story-based QA. However, these works focus on descriptive reasoning rather than procedural reasoning.
II-G Logical queries
@@
Our work is also related to NLP research such as [abdelaziz2021semantic, yang2023llm, zhong2024synthet2c, shah-etal-2024-improving], which employ logical queries to extract information from a KG. This emphasis on curated, domain-specific knowledge is especially valuable for AI assistants operating in contexts where domain expertise, rather than general common-sense reasoning, is crucial for success.
III Procedural Knowledge Graph Construction
@@
Motivated by recent advances in the semi-automated construction of knowledge-based question answering datasets [hoang2024semi], we introduce the Procedural Knowledge Reasoning Question Answering (PKR-QA) dataset. It is built on a Procedural Knowledge Graph (PKG) that integrates information from the COIN training set [tang2019coin], the COIN ontology [tang2019coin], GPT-4o-generated annotations, and external common-sense knowledge from ConceptNet [speer2017conceptnet], followed by human verification. To enable systematic QA generation, we define a set of question templates and associated Cypher queries [francis2018cypher], which are executed over the knowledge graph to retrieve correct answers. In the following sections, we detail the construction of both the knowledge graph and the QA dataset.
III-A Defining PKG’s Schema (PKGS).
@@
A knowledge graph schema provides a high-level abstraction of the graph, specifying the types of entities that exist and the valid relationships that can occur between them. It serves as a blueprint for structuring the data and interpreting the semantics of the graph. In our case, PKGS contains nodes where each node correspond to a entity type ($\mathcal{E}$) and edges represent relation types ($\mathcal{R}$) (Figure 1). The core entity types in PKGS include Domain, Task, Step, Action, Object, Tool, and Purpose. Relation types capture meaningful procedural links, such as task-step associations or tool usage (e.g. HAS_TOOL), and may carry attributes like id, type, or additional semantic metadata. The detailed descriptions and definitions of each entity and relation is given in Section III-C.
III-B Populating PKG.
@@
Based on this schema, we instantiate PKG by populating it with specific entity and relation instances. Each entity $e\in\mathcal{E}$ has three main attributes: type, name, and id. Relations $r\in\mathcal{R}$ represent specific connections between entity instances. To construct this graph, we integrate annotations from the COIN dataset [tang2019coin], fine-grained information extracted using GPT-4o [hurst2024gpt], and commonsense knowledge from ConceptNet [speer2017conceptnet].
@@
COIN-based Data (Domain, Task, Step): We extract the procedural structure from the training split of the COIN dataset, which provides annotations for domains, tasks, and steps. For each unique instance of these entities, we create a node in the graph. We define HAS_TASK edges to link each domain to its associated tasks, and HAS_STEP edges to connect tasks to their constituent steps. To capture procedural flow, we analyze step sequences from all training videos and construct HAS_NEXT_STEP relations that represent the temporal ordering of steps. We also introduce START and END nodes to explicitly mark task boundaries. Each HAS_NEXT_STEP edge is annotated with its observed frequency to model empirical transition likelihoods between steps.
@@
LLM-Augmented Data (Action, Object, Tool): To enhance procedural specificity, we use GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ to extract action-object pairs from step descriptions. For instance, the phrase “remove the tire” yields the action “remove” and the object “tire”. Because COIN lacks explicit tool annotations, we further prompt GPT-4o to infer potential tools for each step. The prompt that we use is shown in Figure 6 (Appendix). The tool lists are manually verified. We standardize terms (e.g., “scissor” $\rightarrow$ “scissors”), consolidate duplicates (e.g., “marker”, “marker pen”), and unify task-specific variants (e.g., “toilet detergent” $\rightarrow$ “detergent”). Some task-sensitive distinctions are preserved (e.g., “filter” $\rightarrow$ “coffee filter”) when relevant to the procedural context.
@@
ConceptNet-based Data (Purpose): In procedural tasks, actions are typically performed—and tools or objects used—with specific purposes. To represent such intent in our graph, we incorporate commonsense knowledge from ConceptNet, a large-scale semantic network that connects words and phrases through meaningful, human-readable relations such as UsedFor, CapableOf, and IsA. Specifically, we extract potential purposes of actions, tools, and objects using the UsedFor and CapableOf relations. To contextualize these purposes within specific procedural steps and tasks, we use GPT-4o to infer task- and step-specific interpretations for each entity’s purpose. This step ensures that the resulting knowledge is not generic but grounded in the procedural context in which the entity is used. (See  Figure 7 – Appendix)
@@
To reduce redundancy and merge semantically similar purposes, we compute pairwise cosine similarities of Sentence-BERT embeddings [reimers-2019-sentence-bert] and apply a similarity threshold of 0.8, validated through manual inspection. Pairs of highly similar purposes are linked using a HAS_SIMILAR_PURPOSE edge. The resulting purpose-augmented graph is stored in Neo4j [^2] https://neo4j.com/, supporting efficient querying and reasoning via Cypher. All contents are manually verified to ensure the reliability of the KG. The constructed KG contains 2,954 unique entities and 12,484 relations. The distributions of entities and relations are shown in Figure 10 of the supplementary.
III-C Detailed definition of PKG
Figure 1(left) shows PKGS, the schema of our procedural knowledge graph, PKG. PKGS consists of 9 nodes (entities, $\mathcal{E}$) and 14 edges (relations, $\mathcal{R}$) connecting these nodes. The nodes represent various types of entities, each capturing a key aspect of procedural knowledge, including:
  • Domain: contains 12 distinct domains which represent the highest-level category or field under which a set of related tasks are grouped.
  • Task: represents a specific activity or goal within a domain that requires a set of coordinated steps to complete. Tasks provide the contextual framework for steps, defining the objective that the steps aim to achieve. It bridge the gap between the high-level domain and the detailed procedural steps.
  • Step: detailing the individual operations necessary to accomplish a specific task. Steps are sequentially organized to capture the logical flow required for successful task completion.
  • Action: represents atomic operation performed as part of a step. Actions are more granular levels that detail how a step is executed. For example, actions such as “scrub”, “fry”, or “pump up” describe the specific motions or manipulations required in each step.
  • Tool: represents a specific tool that can be used to assist in performing a step. These tools are not just generic but are tied to the visual evidence in the procedural context. For example, a “sponge” and “cleaning solution” for the step “scrub the bathtub” under task “clean bathtub” are tools.
  • Object: refers to the item being manipulated or acted upon in a step. Objects are central to the procedural actions as they undergo changes or are used as part of the process. Using previous example, in the step “scrub the bathtub”, the object is the bathtub.
  • Purpose: captures the intention or function behind a specific grounded tool within the context of a task and step. It is important for answers the “what” tool can be used in the step if we want to achieve a certain task. For example, the “sponge” has purpose of “cleaning” under step “scrub the bathtub”. This node type enhances the knowledge graph’s interpretability by explicitly associating tools with their roles in achieving procedural goals.
  • START and END: two dummy nodes that connect to the starting step and ending step, respectively. These two entities are used to identify starting step and ending step of a task.
The edges in PKGS capture interactions, dependencies, and hierarchies between entities. These relations include:
  • HAS_TASK: links a Domain node to its associated Task nodes. It represents the grouping of tasks under a domain.
  • HAS_STEP: connects a Task node to its associated Step nodes, capturing the breakdown of the task.
  • HAS_NEXT_STEP: capture the sequential order between Step nodes within a task. For example, in the task “clean bathtub”, the step “scrub the bathtub” is followed by the step “clean bathtub with water”. This relation helps in modeling the progression of actions.
  • HAS_ACTION: connects a Step node to its corresponding Action nodes, detailing the atomic operations required to execute the step.
  • HAS_TOOL: links a Step node to the Tool nodes that assist in performing the step.
  • HAS_OBJECT: links a Step node to the Object nodes that are manipulated or acted upon during the step.
  • HAS_PURPOSE: connects a GroundedTool node to a Purpose node, capturing the function or intent of the tool in the procedural context.
  • HAS_SIMILAR_PURPOSE: links two Purpose nodes that share similar intents or functions. For instance, the purposes “carry liquids” and “hold liquids” are linked with this relation.
The structure of our PKGS not only captures the hierarchical structure and temporal relationships between various components of procedural knowledge extracted from [tang2019coin], but also extends its expressiveness and utility by incorporating additional layers of information. By enriching the graph with detailed node types and their interconnections, our PKGS provides a more holistic representation of procedural knowledge.
IV Question Answering Dataset Creation